Train A has a speed 20 miles per hour greater than that of train B. If train A travels 230 miles in the same times train B travels 190 miles, what are the speeds of the two trains?

Answer by richwmiller(17219)

t=d/r
230/(b+20)=190/b
b=95
a=115
t=2